A Bibliography of Publications of the USENIX Association: 2010–2019

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

06 February 2024 Version 1.02

Title word cross-reference

3 [San10].

/dev/random [Fer10a, Fer10c, Fer10b, Fer10d, Fer10e, Fer10f, Fer11e, Fer11f, Fer11a, Fer11b, Fer11c, Fer11d, Fer12e, Fer12f, Fer12a, Fer12b, Fer12c, Fer12d].

10 [See10]. 10th [AFF^{+}12]. ’11 [CFK^{+}12, LH11, Gal12]. ’12 [AFF^{+}12, DFF^{+}12]. 14th [ABE^{+}12].

2012 [Gee12b]. 25th [CFK^{+}12].

3 [Bea12f, TAB^{+}10].

5th [HVB12].

A-Wanderin [BE12c]. Accessing [Tan11b]. Accounts [Sch11a]. ACM [HVB12]. Acts [BE10e]. Added [Oet12]. Address [NPFS11]. Administration [BR10, CFK^{+}12, Cou10b, Fis11, Lim10, Sch10]. Adrian [Far12i]. AdSplit [SDW12]. Advanced [Mor11]. Advertising [SDW12]. Age [Her12]. Agile [Fis11]. Ago [Oet12]. Airport [Fer10c]. AIX [Gal10a, Gal11c]. Alchemy [Cou10c]. Alternative [MMEK^{+}10]. Always [BE12d]. American [For12]. Analysis [RBM^{+}10]. Analytics [ZCD^{+}12]. Anatomy [CSJH12]. Android [MV12b]. Anti [RBM^{+}10]. Anti-Virus [RBM^{+}10]. Apache [Shv11]. APIs [BE12d]. Appendix [MV12a]. Appliance [See10]. Applications [MMP^{+}12, SDW12]. Apps [Fer12e]. Architecture [Gal10c, Rev11]. Army [Bac12]. Art [WB12]. Assurances [Cou10b]. Attack [Mud11]. Attacks [CSJH12]. Automatically [HFK^{+}10]. Automating [AF11]. Availability [Cha12]. Available [Ken11].

Back [BWV11]. Backup [Far11a]. Balancing [RRKW12]. Barrelfish [Far10a]. Base [HFK^{+}10]. Based [APVL12]. Basics [Lan11]. Becoming [Bea12a, Fis11]. Beginning [Sim10]. Better [BE11a, UKK^{+}12]. Beyond [Lan11, WW10]. Big [CGZ^{+}12, Dea11]. BIND [See10]. Block [Her12]. Blocks [WC12, WW10]. Bloggers [Dic10]. Board [Cou11, Eva10]. Book [wBC10, CS10, LCS^{+}12, wSS12, ZCS10a, ZCS10b, ZFS10, ZFS11, ZSTF11, ZTM11, ZF11, ZDSG11, ZS11, ZLDC12, ZGSL12, ZL12, ZF12, Z^{+}10]. Books [Dic11]. Both [MV12b, MV12a]. Brain [Jos10c, Jos11f]. Brave [EWS12]. Breaking [Jos10b]. Brute [Zwi10]. Btrfs [Bac12]. Buffer [BLPS11]. Building [Dar11, LB10, See10]. Business [BR10].

Calling [Dic10]. Capabilities [WALK10]. Capsicum [WALK10]. Care [Gal12]. Case [LB10, Mor11, ZTS10]. Cassandra [Far12h]. CDMI [Sli12]. Center [MH11]. Centralized [DH11]. Ceph [MMEK^{+}10]. Cfengine [Bjø10]. Challenges [LH11]. Championships [Kol10]. Changing [Jos11c, Jos12a, Jos12b, Jos12c]. China [Kol10, WC12]. chock [BE12a]. CIEL [MH11]. Client [Ken12a]. Client/Network/Server [Ken12a]. Clones [JWB12]. Cloud [Bur10, Dar11, Gal11a, Gal11b, HKZ^{+}11, Jos11d, Sch10, Wan12]. Clouds [Far12i]. Cluster [LB10, MMP^{+}12]. Cockcroft [Far12i]. Code [JWB12, Sch11b]. Coincidence [HO12, Hug12b]. Collaboration [Mud11]. Collector [Bea12a]. Collide [HO12, Hug12b]. Command [Tan11a, Tan11b]. Command-Line [Tan11a]. Committee [Eva10]. Commodity [LB10]. Common [Fer10d]. Compared [Bjø10]. Comparing [Gal10a, Gal11c]. Competition [Dea11]. Compliance [Gee10]. Components [LB10]. Computation [BLPS11]. Computer [HVB12]. Computers [MV12b, MV12a]. Computing [HFK^{+}10]. ConfAid [AF11]. Conference [AFF^{+}12, Gee12b, CFK^{+}12]. Configuration [AF11, Cou10a, Lue11]. Consistent [EWS12]. Consumption [Bjø10]. Contenders [San10]. Contents [Edi12a, Edi12b]. Continued [Hug12b]. Contracts [PK10]. Create [UKK^{+}12]. Criticality [Gee12a]. Crom [Jos11e]. Crossbow [Gal10b]. Crosstalk [Jos12d]. CSV [BE12c]. Current [TAB^{+}10]. Cybersecurity [For12].

D [San10]. Dad [Ebe12]. Dancer [BE11c]. Danger [Sch11b, BE12f]. Dark [Fer11f, Fer12f, Har12]. Data [Bea12b, Cha12, CGZ^{+}12, EWS12, Far12a, Hug12a, MH11, ZCD^{+}12]. Databases [Tan11b]. Date [BE10f]. Days [BE12d]. DBURLs [Tan11b]. Debatably [Fer10b]. Decentralized [DH11]. Deduplication [Gal10d]. Delivery [Gee12b]. Design [Khu12]. Desktop [MV12b, MV12a]. Desktops [MWHW11]. Detecting [APVL12, HFK^{+}10]. dev [Fer10a, Fer10c, Fer10b, Fer10d, Fer10e, Fer10f, Fer11e, Fer11f, Fer11a, Fer11b, Fer11c, Fer11d, Fer12e, Fer12f, Fer12a, Fer12b, Fer12c, Fer12d]. Devices [Kir10, vD10a]. DevOps [Deb11, LH11]. DFS [JBFL10]. DGA [APVL12]. DGA-Based [APVL12]. Different [Tan11b]. Directors [Cou11, Eva10]. Directory [NPFS11]. Disaster [Fer10b]. Disasters [HO12, Hug12b]. Distributed [Cha12, MMEK^{+}10]. Distribution [RBM^{+}10]. Distributions [JWB12]. DNS [See10, Vix11]. Do [BE11a, Fis11, Gut11, Jos10b]. Does [BE10d]. Don’t [CSR10]. Doug [Far11a]. Down [BE12d, Ken12b]. Draft [Gee12b]. Durability [Cha12]. Dynamics [Far11a].

Easier [Fis11]. Editor [BE11b]. Effect [RBM^{+}10]. Election [Cou11, Eva10]. Email [Kir10]. Embedded [vD10b]. Enterprise [Gal11a, Gal11b, Gal10a, Gal11c]. Entire [JWB12]. Errors [Het11]. Exadata [Gal10c]. Exchange [McK10]. Exploit [BLPS11]. Express [Gal11d]. Extended [Kol10]. Extending [Wan12]. Extensions [BWV11]. Extreme [KCT12].

Fake [RBM^{+}10]. Fall [Har12]. Family [BE10a]. Fanged [HO12, Hug12b]. Farmer [Far12a]. Fashioned [WB12]. FAST [AFF^{+}12, ZCD^{+}12]. Fault [EWS12]. Fault-tolerant [EWS12]. Features [Gal11c]. Fighting [ZTS10]. File [AFF^{+}12, Cha12, Dar11, JBFL10, MMEK^{+}10, ZTS10]. Files [BE12b, CSR10, WW10]. Finding [Hug12a, JWB12]. Fire [Jos12f]. Firefox [MV12b]. Firewall [WC12]. First [BE10f]. Fit [Wan12]. Five [Fer10d]. Flash [JBFL10]. Flexible [HKZ^{+}11]. Force [Zwi10]. Forget [BE10f]. Free [Gal10d]. Functionality [RRKW12]. Future [BWV11, Gee12a, Gee12b].

Galvin [Gal11a, Gal11b]. Game [Jos11c, Jos12a, Jos12b, Jos12c]. Ganglia [Jos10c, Jos11f]. Gazing [Jos11d]. Gems [Jos11a]. Get [BE10b, BE11c, BE12d]. Getting [Lue11]. Gift [Jos12f]. Give [BE11c]. Glasnost [Her12]. Glass [Lim10]. GNU [Tan11a, Tan11b]. Go [BE12c]. Good [BE11c, WB12]. Got [BE11d]. Great [WC12]. Growth [Shv10]. Guesses [Hug12a]. Guide [Fer10b].

H [BE11e]. H-TTP [BE11e]. Hadoop [Cha12, LB10, MMEK^{+}10, Shv11, ZCD^{+}12]. Hadoop/MapReduce [LB10]. Halting [SPBS11]. Handful [Jos11a]. Hard [Fis11, Het11, Jos10b, WC12]. Hardware [HFK^{+}10]. HBase [Khu12]. HDFS [Shv10]. Headless [Sch11a]. Heads [Far12i]. Help [BE10b]. Helpful [Fer10b]. Helping [UKK^{+}12]. High [ABE^{+}12, EWS12, Kol10]. Highly [Ken11]. Hither [BE11f]. Hosted [McK10]. Hosts [Ken11]. HPTS [ABE^{+}12]. HTTPS [Lan11]. Hughes [Far11a]. Human [HVB12]. HyperDex [EWS12]. Hypervisors [Rev11].

Identifier [Pis10]. Ignorance [Zwi10].

Impacts [Het11]. Implementation [Rev11]. Implications [CGZ^{+}12]. Import [Bea12c]. Improve [For12]. Improving [PK10]. Incast [CGZ^{+}12]. Information [Fer10d, Gee12b]. Informed [MMP^{+}12]. Insecurity [SPBS11]. Insights [Rev11]. Installation [CFK^{+}12]. Installing [LB10]. Integrity [Hug12a]. Interaction [HVB12]. Interactive [ZCD^{+}12]. International [ABE^{+}12, Dea11]. Internet [Ebe12, NPFS11, Pis10]. Interview [Far10a, Far11a, Far11b, Far12a, Far12i, Far12h]. Introducing [WALK10]. Introduction [EWS12, Khu12, Sli12]. IPv6 [BWV11, Ebe12, Rya10]. iVoyeur [Jos10b, Jos10c, Jos10d, Jos10e, Jos10f, Jos10a, Jos11b, Jos11e, Jos11d, Jos11c, Jos11a, Jos11f, Jos12d, Jos12a, Jos12b, Jos12c, Jos12f, Jos12e].

Jacob [Far12a]. Journey [Sim10].

Kernel [KPAK12]. Keys [Gut11]. kGuard [KPAK12]. Kindness [BE10e]. Knife [Bac12]. Know [Rya10].

Large [CFK^{+}12]. Last [KCT12].

Last-Millennium [KCT12]. Latest [Dic11]. Law [KS10]. LEET [DFF^{+}12]. Less [BE12e, Fer10e]. Let [BE10b]. Letting [Fis11]. Lewis [KS10]. Lies [Hug12a].

Lightweight [KPAK12]. Limits [Shv10]. Line [Tan11a, Tan11b]. LISA [CFK^{+}12, Kam10, LH11, Sim10]. Live [PK10]. Logging [DH11]. Logs [Dil10]. Look [BE11a, Ker10]. Looking [For12].

Machines [Fis11, Rev11, BLPS11]. makacheesa [BE12a]. Making [BE10c, Dil10, Fis11]. Mala [BE12a]. Malicious [HFK^{+}10]. Malware [APVL12]. Man [BE10a]. Management [Cou10a, Fer10f, HVB12, Lue11, MWHW11]. Managing [Far12a]. Many [HO12, Hug12b]. Many-Fanged [HO12, Hug12b]. MapReduce [LB10]. Master [Bea12a]. Matter [BE10d]. Matters [Gal10c]. May [Gee12b]. Me [BE10b, BE11e, BE12d]. Mean [BE12b]. Means [BE11e]. Memoriam [KS10]. Mesos [HKZ^{+}11]. Message [Her12]. Microsoft [Her12]. Migrating [McK10]. Milford [Far12h]. Millennium [KCT12]. Mind [BE12b]. Minder [Bur10]. MINIX [TAB^{+}10]. Misconceptions [Fer10d]. Mobile [Kir10, MV12b, MV12a, WB12]. Model [LRA10]. Module [Bea12e]. Mojolicious [BE11d]. Molecular [Far11a]. Mosh [WB12]. Mountain [Gee12b]. Multikernel [Far10a]. Multipath [BHR12]. Multiprocessing [Bea12e]. Musings [Far10b, Far10c, Far10d, Far10e, Far10f, Far10g, Far11c, Far11d, Far11e, Far11f, Far11g, Far11h, Far12b, Far12c, Far12d, Far12e, Far12f, Far12g]. My [BE11c, BE11d]. Myths [Fer10c].

Nagios [Jos12e]. Name [NPFS11]. Namespace [MWHW11]. NAS [Gal10e]. Nathan [Far12h]. Need [Rya10]. Netflix [Far12i]. Network [Fer10e, Ken12a, Mud11, SPBS11, She11]. Networking [Wan12]. Never [BE10f]. NFSv4 [McD12]. Nocebo [RBM^{+}10]. Nominal [Gee12b]. Nominating [Eva10].

NoSQL [Bur11]. Notes [HD12]. Noticing [Oet12]. NSDI [DFF^{+}12].

o [Jos10d, Jos10e, Jos10f]. Obesity [ZTS10]. Object [BE10f]. ObPwd [MV12b]. Old [WB12]. Old-Fashioned [WB12]. Open [Gal10d]. Other [Fer10c, Vix11]. Our [You10a, You10b]. Overcoming [HFK^{+}10]. Overflows [BLPS11]. Overview [BHR12].

P2P [PK10]. Packets

[Jos10d, Jos10e, Jos10f]. Pandas [Bea12b]. Parable [Lim10]. Parallel [Tan11a]. Parallelism [KCT12]. Parenthetically [Fer11e]. Part [Gal11b, Jos10d, Jos10e, Jos10f, Jos12a, Jos12b, Jos12c]. Party [RRKW12]. Passwords [MV12b, MV12a, UKK^{+}12]. Patience [HO12, Hug12b]. Payment [Far11b]. Peculiarities [vD10a]. pee [BE12a]. Performance [ABE^{+}12, EWS12, PK10]. Perhaps [BE10d]. Perl [BE10f, BE10b, BE10c, BE10e, BE10d, BE10a, BE11f, BE11e, BE11c, BE11d, BE11a, BE11b, BE12c, BE12f, BE12d, BE12e, BE12b, BE12a, Oet12]. Persistent [Mor11]. Perspective [Deb11]. Pete [Gal10b, Gal10d, Gal10c, Gal10f, Gal10e, Gal10a, Gal11c, Gal11d]. Phenomenology [Cou10a]. Pie [Ken12a]. Pinpoint [WC12]. Planning [Fer10b]. Platform [LRA10]. Pockets [Jos10d, Jos10e, Jos10f]. Pockets-o-Packets [Jos10d, Jos10e, Jos10f]. Popping [Pol12]. Power [Tan11a, Pol12]. Practical [BE10f, BE10b, BE10c, BE10e, BE10d, BE10a, BE11f, BE11e, BE11c, BE11d, BE11a, BE11b, BE12c, BE12f, BE12d, BE12e, BE12b, BE12a, WALK10]. Primer [Fer10f]. Privacy [BWV11, RRKW12]. ‘Problem’ [Gal10e]. Problems [SPBS11]. Processing [Bea12b]. Production [She11]. Products [Gal10f]. Profile [BE11a]. Programming [BLPS11, Cou10c, Dea11, Gar12, Kol10, MH11].

Project [Fer10f]. Protection [KPAK12]. Protocol [Ebe12]. Protocols [Fer10e]. Proven [LH11]. Proverb [Fer11f]. Provisioning [MMP^{+}12]. Puppet [Bjø10]. Python [Bea12f].

Question [Jos10a].

‘R’ [Bea12d]. Radio [vD10a]. Rainy [BE12d]. Random [BE10e, Fer10a, Fer10c, Fer10b, Fer10d, Fer10e, Fer10f, Fer11e, Fer11f, Fer11a, Fer11b, Fer11c, Fer11d, Fer12e, Fer12f, Fer12a, Fer12b, Fer12c, Fer12d]. Really [BE10d]. ReDeBug [JWB12]. Redefining [Cou10b]. RedHat [Gal10a, Gal11c]. Regular [BE10b]. Rejectionists [Gee12a]. Removing [HFK^{+}10]. Replacement [Bea12d]. Report [Eva10, HVB12, TAB^{+}10]. Reports [CFK^{+}12]. Research [Far12a, TAB^{+}10]. Resource [Bjø10, HKZ^{+}11]. RESTful [NPFS11]. Restricted [Sch11a]. Reviews [wBC10, CS10, LCS^{+}12, wSS12, ZCS10a, ZCS10b, ZFS10, ZFS11, ZSTF11, ZTM11, ZF11, ZDSG11, ZS11, ZLDC12, ZGSL12, ZL12, ZF12, Z^{+}10]. Revisiting [BWV11]. Rhetoric [Fer12f]. Rise [APVL12, Har12]. Risk [Gee12a]. Ritual [Cou10c]. Robinson [BE12f]. Robustness [PK10]. Rocky [Gee12b]. Roscoe [Far10a].

Safely [She11]. Saps [Fer12e]. Savage [Far11b]. Scalability [Shv10, Shv11]. Scalable [EWS12, MMEK^{+}10]. Scale [Jos10a]. scc [MMP^{+}12]. Schema [Khu12]. School [Kol10]. Secrets [Bea12e]. Secure [Kir10, Vix11]. Securing [Pis10]. Security [Fer10c, Fer10d, Gee12a, Gee12b, LRA10]. Seen [Dic11]. Sense [Dil10]. Separating [SDW12]. Server [Ken12a, Her12]. Service [NPFS11]. Services [NPFS11]. Serving [Lan11]. ShareMeNot [RRKW12]. Sharing [HKZ^{+}11, Mud11]. Shell [WB12, Mud11]. Shells [Sch11a]. Short [Dic11]. Should [Gal12]. Side [Ken11]. Silicon [Har12]. SIP [CSJH12]. Size [BE10d]. Skywriting [MH11]. Small [vD10b]. Smartphone [SDW12]. Snaps [Kol10]. Social [RRKW12]. Solaris [Gal10a, Gal11c, Gal11d, Gal12, See10]. Sooooo [KCT12]. Source [Gal10d]. Spam [Far11b]. Spark [ZCD^{+}12]. Speaking [Fer11e]. Specialized [ZTS10]. Spreadsheet [BE12c]. SQL [Tan11b]. SSH [Gut11]. SSR [Pis10]. Stables [BE12b]. Stack [SPBS11]. Started [Lue11]. State [Gal11a, Gal11b, WB12]. State-of-the-Art [WB12]. Status [Gal10f, TAB^{+}10]. Stefan [Far11b]. Storage [AFF^{+}12, Bac12, Het11, JBFL10, MMP^{+}12]. Stores [EWS12]. Strangers [CSR10]. Strategies [Far11a]. Streak [Kol10]. Streaming [PK10]. Stress [HO12, Hug12b]. Study [LB10, Mor11]. Stuff [BE10c]. Successful [Fer10e]. Summer [Pol12]. Sun [Gal10b, Gal10d, Gal10c, Gal10f, Gal10e, Gal10a, Gal11c, Gal11d]. Supplies [Pol12]. Support [Oet12]. Swiss [Bac12]. Symbian [LRA10]. Symposium [HVB12]. Sysadmin [Deb11, Dic11, Fis11]. System [BR10, CFK^{+}12, Cou10b, Dar11, Fis11, Het11, JBFL10, Lim10, MMEK^{+}10, Sch10, ZTS10, Cha12]. Systems [ABE^{+}12, Pis10, ZTS10, vD10b].

Table [Edi12a, Edi12b]. Take [CSR10]. Takes [Ken12b]. Taking [BE12e]. Tale [HO12, Hug12b]. Taming [CB10]. Tasks [Cou10b]. Tasting [Ken12a]. TCP [BHR12, CGZ^{+}12]. Teaching [Sch10]. Team [Dea11]. Technological [Cou10c]. Technologies [AFF^{+}12]. Testability [Het11]. Testbed [She11]. Testing [Ken11]. Thanks [You10a, You10b]. Theme [LH11]. Theory [BLPS11]. Things [Gal10b, Gal10d, Gal10c, Gal10f, Gal10e, Gal10a, Gal11a, Gal11b, Gal11c, Gal11d].

Third [RRKW12]. Third-Party [RRKW12]. Those [BE12b]. Threats [Mor11]. Three [Bea12f]. Time [Bjø10]. Timothy [Far10a]. Tiny [BE11c]. Together [Jos11b]. Tolerance [Gee12a]. tolerant [EWS12]. Tool [Tan11a, Tan11b]. Tools [BE10f, BE10b, BE10c, BE10e, BE10d, BE10a, BE11f, BE11e, BE11c, BE11d, BE11a, BE11b, BE12c, BE12f, BE12d, BE12e, BE12b, BE12a]. Topics [Dic11]. Tor [WC12]. Torrent [CB10]. Tracking [APVL12]. Trail [Far11b]. Transaction [ABE^{+}12]. Transit [Rya10]. Transport [Ken11]. Traveled [BE12e]. Trends [Het11]. trom [BE12a]. Troubleshooting [AF11]. Truth [Hug12a]. TTP [BE11e].

Understanding [CGZ^{+}12, Mor11]. Undocumented [BE12d]. Unicode [Gar12, Oet12]. UNIX [WALK10]. Unpatched [JWB12]. Unrandomized [Sch11b]. Untrusted [HFK^{+}10]. Update [Shv11]. USA [Dea11, Kol10]. USENIX [AFF^{+}12, CFK^{+}12, Cou11, Dic11, Eva10]. User [Sch11a]. Users [Gut11, UKK^{+}12]. Uses [Vix11]. Using [See10, She11, Tan11b].

V2 [Gal10c]. Value [BR10, LH11]. Verify [Gut11]. Virtual [MWHW11, Rev11, See10]. Virtualization [Fer11f, Gal11c]. Virtualized [JBFL10]. Virus [RBM^{+}10]. VMware [See10]. VoIP [Ker10]. Volunteers [You10a, You10b]. Vulnerabilities [Ker10]. Vulnerable [Gee10]. VXLAN [Wan12].

Waitress [Lim10]. Wanderin [BE12c]. Warning [BE12f, BE12f]. Water [Lim10]. Web [NPFS11, RBM^{+}10, San10]. ‘Weird [BLPS11]. Whither [Kam10]. Widgets [RRKW12]. Will [BE12f]. Winning [Kol10]. Wins [Dea11, Kol10]. Within [For12]. Without [Oet12]. Woe [HO12, Hug12b].

Work [Fis11]. Working [BE11d]. Workloads [CGZ^{+}12]. Workshop [ABE^{+}12]. World [DH11, EWS12, Hug12a, Kol10].

x86 [Rev11]. XI [Jos12e]. XPath [BE12e].

Years [Bea12f, Oet12]. Yon [BE11f].

References

Armbrust:2012:IWH

[ABE^{+}12] Michael Armbrust, Yingyi Bu, Aaron Elmore, Rik Farrow, Eugenia Gabrielova, Hatem Mahmoud, Andy Pavlo, Steve Revilak, and Pinar Tozun. 14th International Workshop on High Performance Transaction Systems (HPTS). ;login: the USE-NIX Association newsletter, 37 (1):??, February 2012. CODEN LOGNEM. ISSN 1044-6397.

                                                   URL https://www.usenix.
                                                   org/publications/login/february-
                                                   2012/14th-international-workshop-
                                                   high-performance-transaction-
                                                   systems.

Attariyan:2011:ACT

[AF11] Mona Attariyan and Jason Flinn. Automating configuration troubleshooting with ConfAid. ;login: the USENIX Association newsletter, 36(1): ??, February 2011. CODEN LOGNEM. ISSN 1044-6397.

                                                   URL https://www.usenix.
                                                   org/publications/login/february-
                                                   2011-volume-36-number-1/automating-
                                                   configuration-troubleshooting-
                                                   confaid.

Arteaga:2012:UCF

[AFF^{+}12] Dulcardo Arteaga, Rik Farrow, Daniel Fryer, Doowon Kim, Michelle L. Mazurek, Dutch Meyer, Swapnil Patil, and Yiqi Xu. 10th USE-NIX Conference on File and Storage Technologies (FAST ’12). ;login: the USENIX Association newsletter, 37(3): ??, June 2012. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/june-
2012-volume-37-number-3/10th-
usenix-conference-file-and-
storage-technologies.

Antonakakis:2012:DTR

[APVL12] Manos Antonakakis, Roberto Perdisci, Nikolaos Vasiloglou, and Wenke Lee. Detecting and tracking the rise of DGA-based malware. ;login: the USENIX Association newsletter, 37(6): ??, December 2012. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/december-
2012-volume-37-number-6/detecting-
and-tracking-rise-dga-based-
malware.

Bacik:2012:BSA

[Bac12]  Josef Bacik.  Btrfs: The

Swiss Army knife of storage. ;login: the USENIX Association newsletter, 37(1):7– 15, February 2012. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/february-

                                        2012/btrfs-swiss-army-knife-
                                        storage.

Blank-Edelman:2010:PPTf

[BE10a] David N. Blank-Edelman. Practical Perl tools: Family man. ;login: the USENIX Association newsletter, 35(6):??, December 2010. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/december-
                                        2010-volume-35-number-6/practical-
                                        perl-tools-family-man.

Blank-Edelman:2010:PPTb

[BE10b] David N. Blank-Edelman. Practical Perl tools: Let me help you get regular. ;login: the USE-NIX Association newsletter, 35 (2):??, April 2010. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/april-
                                        2010-volume-35-number-2/practical-
                                        perl-tools-let-me-help-you-
                                        get-regular.

Blank-Edelman:2010:PPTc

[BE10c] David N. Blank-Edelman. Practical Perl tools: Making stuff up. ;login: the USENIX Association newsletter, 35(3): ??, June 2010. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/june-
                                        2010-volume-35-number-3/practical-
                                        perl-tools-making-stuff.

Blank-Edelman:2010:PPTe

[BE10d] David N. Blank-Edelman. Practical Perl tools: Perhaps size really does matter. ;login: the USENIX Association newsletter, 35(5):??, October 2010. CO-DEN LOGNEM. ISSN 1044-

6397.  URL https://www.
usenix.org/publications/login/
october-2010-volume-35-number-
5/practical-perl-tools-perhaps-
size-really-does.

Blank-Edelman:2010:PPTd

[BE10e] David N. Blank-Edelman. Practical Perl tools: Random acts of kindness. ;login: the USE-NIX Association newsletter, 35 (4):??, August 2010. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/august-
2010-volume-35-number-4/practical-
perl-tools-random-acts-kindness.

Blank-Edelman:2010:PPTa

[BE10f] David N. Blank-Edelman. Practical Perl tools: You never forget your first date (object). ;login: the USENIX Association newsletter, 35(1): ??, February 2010. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/february-
2010-volume-35-number-1/practical-
perl-tools-you-never-forget-
your-first.

Blank-Edelman:2011:PPTe

[BE11a] David N. Blank-Edelman. Practical Perl tools: Do I look better in profile? ;login: the USE-NIX Association newsletter, 36 (5):??, October 2011. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/october-
                                        2011-volume-36-number-5/practical-
                                        perl-tools-do-i-look-better-
                                        profile.

Blank-Edelman:2011:PPTf

[BE11b] David N. Blank-Edelman. Practical Perl tools: From the Editor. ;login: the USENIX Association newsletter, 36(6): ??, December 2011. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/december-
                                        2011-volume-36-number-6/practical-
                                        perl-tools-editor.

Blank-Edelman:2011:PPTc

[BE11c] David N. Blank-Edelman. Practical Perl tools: Give as good as you get, my tiny dancer. ;login: the USE-NIX Association newsletter, 36 (3):??, June 2011. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/june-
                                        2011-volume-36-number-3/practical-
                                        perl-tools-give-good-you-
                                        get-my-tiny.

Blank-Edelman:2011:PPTd

[BE11d] David N. Blank-Edelman. Practical Perl tools: Got my mojolicious working. ;login: the USE-NIX Association newsletter, 36 (4):??, August 2011. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/august-
                                        2011-volume-36-number-4/practical-

perl-tools-got-my-mojolicious-
working.

Blank-Edelman:2011:PPTb

[BE11e] David N. Blank-Edelman. Practical Perl tools: H-T… TP

— that’s what it means to me. ;login: the USENIX Association newsletter, 36(2): ??, April 2011. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/april-
2011-volume-36-number-2/practical-
perl-tools-h-t-tp%E2%80%94thats-
what-it-means-me.

Blank-Edelman:2011:PPTa

[BE11f] David N. Blank-Edelman. Practical Perl tools: Hither and yon. ;login: the USENIX Association newsletter, 36(1): ??, February 2011. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/february-
2011-volume-36-number-1/practical-
perl-tools-hither-and-yon.

Blank-Edelman:2012:PPTf

[BE12a] David Blank-Edelman. Practical Perl tools: ‘mala trom pee chock makacheesa’. ;login: the USENIX Association newsletter, 37(6):??, December 2012. CODEN LOGNEM. ISSN 1044-6397. URL https://www.

usenix.org/publications/login/
december-2012-volume-37-number-
6/practical-perl-tools-%E2%
80%9Cmala-trom-pee-chock.

Blank-Edelman:2012:PPTe

[BE12b] David Blank-Edelman. Practical Perl tools: Mind those stables, I mean files. ;login: the USENIX Association newsletter, 37(5):??, October 2012. CODEN LOGNEM. ISSN 1044-6397. URL https://www.

                                        usenix.org/publications/login/
                                        october-2012-volume-37-number-
                                        5/practical-perl-tools-mind-
                                        those-stables-i-mean.

Blank-Edelman:2012:PPTa

[BE12c] David N. Blank-Edelman. Practical Perl tools: CSV and the spreadsheet go A-wanderin’. ;login: the USENIX Association newsletter, 37(1):??, February 2012. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/february-
                                        2012/practical-perl-tools-
                                        csv-and-spreadsheet-go-wanderin%
                                        E2%80%99.

Blank-Edelman:2012:PPTc

[BE12d] David N. Blank-Edelman. Practical Perl tools: Rainy days and undocumented APIs always get me down. ;login: the USE-NIX Association newsletter, 37 (3):??, June 2012. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/june-
                                        2012-volume-37-number-3/practical-
                                        perl-tools-rainy-days-and-
                                        undocumented.

Blank-Edelman:2012:PPTd

[BE12e] David N. Blank-Edelman. Practical Perl tools: Taking the XPath less traveled. ;login: the USENIX Association newsletter, 37(4):??, August 2012. CO-DEN LOGNEM. ISSN 1044-

6397.  URL https://www.
usenix.org/publications/login/
august-2012-volume-37-number-
4/practical-perl-tools-taking-
xpath-less-traveled.

Blank-Edelman:2012:PPTb

[BE12f] David N. Blank-Edelman. Practical Perl tools: Warning! Warning! Danger, Will Robinson! ;login: the USE-NIX Association newsletter, 37 (2):??, April 2012. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/april-
2012/practical-perl-tools-
warning-warning-danger-will-
robinson.

Beazley:2012:BMC

[Bea12a] David Beazley. Becoming a Master collector. ;login: the USE-NIX Association newsletter, 37 (3):??, June 2012. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/june-
2012-volume-37-number-3/becoming-
master-collector.

Beazley:2012:DPP

[Bea12b] David Beazley. Data processing with Pandas. ;login: the USE-NIX Association newsletter, 37 (6):??, December 2012. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/december-

                                        2012-volume-37-number-6/data-
                                        processing-pandas.

Beazley:2012:I

[Bea12c] David Beazley. Import that! ;login: the USENIX Association newsletter, 37(4): ??, August 2012. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/august-
                                        2012-volume-37-number-4/import.

Beazley:2012:RR

[Bea12d] David Beazley. ‘R’ is for replacement. ;login: the USE-NIX Association newsletter, 37 (2):??, April 2012. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/april-
                                        2012/%E2%80%9Cr%E2%80%9D-
                                        replacement.

Beazley:2012:SMM

[Bea12e]  David Beazley.  Secrets of

the multiprocessing module. ;login: the USENIX Association newsletter, 37(5):??, October 2012. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/october-
                                        2012-volume-37-number-5/secrets-
                                        multiprocessing-module.

Beazley:2012:TYP

[Bea12f] David Beazley. Three years of Python 3. ;login: the USE-NIX Association newsletter, 37 (1):??, February 2012. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/february-
2012/three-years-python-3.

Bonaventure:2012:OMT

[BHR12] Olivier Bonaventure, Mark Handley, and Costin Raiciu. An overview of multipath TCP. ;login: the USENIX Association newsletter, 37(5): ??, October 2012. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/october-
2012-volume-37-number-5/overview-
multipath-tcp.

Bjorgeengen:2010:PCC

[Bjø10]  Jarle Bjørgeengen.  Pup-

pet and Cfengine compared: Time and resource consumption. ;login: the USENIX Association newsletter, 35(1): ??, February 2010. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/february-
2010-volume-35-number-1/puppet-
and-cfengine-compared-time-
and-resource.

Bratus:2011:EPB

[BLPS11] Sergey Bratus, Michael E. Locasto, Len Sassaman Meredith L. Patterson, and Anna Shubina. Exploit programming: From buffer overflows to ‘Weird Machines’ and theory of computation. ;login: the USENIX Association newsletter, 36(6):13– 21, December 2011. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.

                                        org/publications/login/december-
                                        2011-volume-36-number-6/exploit-
                                        programming-buffer-overflows-
                                        weird.

Burgess:2010:BVS

[BR10] Mark Burgess and Carolyn Rowland. The business value of system administration. ;login: the USENIX Association newsletter, 35(2):??, April 2010. CO-DEN LOGNEM. ISSN 1044-

                                        6397.  URL https://www.
                                        usenix.org/publications/login/
                                        april-2010-volume-35-number-
                                        2/business-value-system-administration.

Burgess:2010:CM

[Bur10] Mark Burgess. ‘the Cloud minders’. ;login: the USENIX Association newsletter, 35(1): ??, February 2010. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/february-
                                        2010-volume-35-number-1/cloud-
                                        minders.

Burd:2011:N

[Bur11] Greg Burd. NoSQL. ;login: the USENIX Association newsletter, 36(5):??, October 2011. CODEN LOGNEM. ISSN 1044-6397. URL https://www.

                                        usenix.org/publications/login/
                                        october-2011-volume-36-number-
                                        5/nosql.

Barrera:2011:BFR

[BWV11] David Barrera, Glenn Wurster, and P. C. Van Oorschot. Back to the future: Revisiting IPv6 privacy extensions. ;login: the USENIX Association newsletter, 36(1): ??, February 2011. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/february-
2011-volume-36-number-1/back-
future-revisiting-ipv6-privacy-
extensions.

Choffnes:2010:TT

[CB10] David Choffnes and Fabián E.

Bustamante. Taming the Torrent. ;login: the USENIX Association newsletter, 35(1): ??, February 2010. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/february-
2010-volume-35-number-1/taming-
torrent.

Chow:2012:RUL

[CFK^{+}12] Ming Chow, Rik Farrow, David Klann, Erinn Looney-Triggs, Cory Lueninghoener, Scott Murphy, Timothy Nelson, Thang Nguyen, Carolyn Rowland, Josh Simon, Deborah Wazir, and Michael Wei. Reports from USENIX LISA ’11: 25th Large Installation System Administration Conference. ;login: the USE-NIX Association newsletter, 37 (2):??, April 2012. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/april-
2012/reports-usenix-lisa-%
E2%80%9911-25th-large-installation-
system-administration.

Chen:2012:UTI

[CGZ^{+}12] Yanpei Chen, Rean Griffith, David Zats, Anthony D. Joseph, and Randy Katz. Understanding TCP incast and its implications for big data workloads. ;login: the USENIX Association newsletter, 37(3): ??, June 2012. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/june-
                                        2012/understanding-tcp-incast.

Chansler:2012:DAD

[Cha12] Robert J. Chansler. Data availability and durability with the Hadoop Distributed File System. ;login: the USENIX Association newsletter, 37(1): ??, February 2012. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/february-
                                        2012/data-availability-and-
                                        durability-hadoop-distributed-
                                        file-system.

Couch:2010:CMP

[Cou10a]  Alva Couch.  Configura-

tion management phenomenology. ;login: the USENIX Association newsletter, 35(1): ??, February 2010. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/february-
                                        2010-volume-35-number-1/configuration-
                                        management-phenomenology.

Couch:2010:TAR

[Cou10b] Alva L. Couch. From tasks to assurances: Redefining system administration. ;login: the USE-NIX Association newsletter, 35 (2):??, April 2010. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/april-
2010-volume-35-number-2/tasks-
assurances-redefining-system-
administration.

Couch:2010:PTR

[Cou10c] Alva L. Couch. Programming with technological ritual and alchemy. ;login: the USE-NIX Association newsletter, 35 (3):??, June 2010. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/june-
2010-volume-35-number-3/programming-
technological-ritual-and-
alchemy.

Couch:2011:EUB

[Cou11] Alva L. Couch. 2012 election for the USENIX board of directors. ;login: the USENIX Association newsletter, 36(6): ??, December 2011. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/december-
2011-volume-36-number-6/2012-
election-usenix-board-directors.

Ching:2010:BRb

[CS10] Brandon Ching and Sam Stover.

Book reviews. ;login: the USE-NIX Association newsletter, 35 (3):??, June 2010. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/june-
2010-volume-35-number-3/book-
reviews.

Ceron:2012:ASA

[CSJH12] João Marcelo Ceron, Klaus Steding-Jessen, and Cristine Hoepers. Anatomy of SIP attacks. ;login: the USENIX Association newsletter, 37(6): ??, December 2012. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/december-
                                        2012-volume-37-number-6/anatomy-
                                        sip-attacks.

Checkoway:2010:DTL

[CSR10] Stephen Checkoway, Hovav Shacham, and Eric Rescorla. Don’t take LaTEX files from strangers. ;login: the USENIX Association newsletter, 35(4): 17–22, August 2010. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/august-
                                        2010-volume-35-number-4/dont-
                                        take-latex-files-strangers.

Darcy:2011:BCF

[Dar11] Jeff Darcy. Building a Cloud file system. ;login: the USE-NIX Association newsletter, 36 (3):??, June 2011. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/june-
                                        2011-volume-36-number-3/building-
                                        cloud-file-system.

Dean:2011:UTW

[Dea11] Brian Dean. USA team wins big at international programming competition. ;login: the USE-NIX Association newsletter, 36 (6):??, December 2011. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/december-
2011-volume-36-number-6/usa-
team-wins-big-international-
programming.

Debois:2011:DSP

[Deb11]  Patrick Debois.  DevOps

from a sysadmin perspective. ;login: the USENIX Association newsletter, 36(6):??, December 2011. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/december-
2011-volume-36-number-6/devops-
sysadmin-perspective.

Dixit:2012:NL

[DFF^{+}12] Advait Abhay Dixit, Rik Farrow, Andrew Ferguson, Katrina LaCurts, Marcelo Martins, Karthik Nagaraj, Kevin Ngo, Will Scott, Manuel Egele, and Engin Kirda. NSDI ’12 and LEET ’12. ;login: the USE-NIX Association newsletter, 37 (4):??, August 2012. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/august-
2012-volume-37-number-4/9th-
usenix-symposium-networked-
systems-design-and.

Donn:2011:CLD

[DH11] James Donn and Tim Hartmann. Centralized logging in a decentralized world. ;login: the USENIX Association newsletter, 36(2):??, April 2011. CO-DEN LOGNEM. ISSN 1044-

                                        6397.  URL https://www.
                                        usenix.org/publications/login/
                                        april-2011-volume-36-number-
                                        2/centralized-logging-decentralized-
                                        world.

Dickison:2010:CAB

[Dic10] Anne Dickison. Calling all bloggers. ;login: the USE-NIX Association newsletter, 35 (2):??, April 2010. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/april-
                                        2010-volume-35-number-2/calling-
                                        all-bloggers.

Dickison:2011:YSL

[Dic11]  Anne Dickison.  Have you

seen the latest USENIX short topics in sysadmin books? ;login: the USENIX Association newsletter, 36(1):??, February 2011. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/february-
                                        2011-volume-36-number-1/have-
                                        you-seen-latest-usenix-short-
                                        topics.

Dilley:2010:MSL

[Dil10] Ron Dilley. Making sense of logs. ;login: the USENIX Association newsletter, 35(4): ??, August 2010. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/august-
                                        2010-volume-35-number-4/making-
                                        sense-logs.

Ebersman:2012:IYD

[Ebe12] Paul Ebersman. IPv6: It’s not your Dad’s Internet Protocol. ;login: the USENIX Association newsletter, 37(2): ??, April 2012. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/april-
2012/ipv6-it%E2%80%99s-not-
your-dad%E2%80%99s-internet-
protocol.

Editor:2012:TCa

[Edi12a]  Editor.  Table of contents.

;login: the USENIX Association newsletter, 37(4): ??, August 2012. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/august-
2012-volume-37-number-4/table-
contents.

Editor:2012:TCb

[Edi12b]  Editor.  Table of contents.

;login: the USENIX Association newsletter, 37(5):??, October 2012. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/october-
2012-volume-37-number-5/table-
contents.

Evard:2010:NCR

[Eva10] Rémy Evard. Nominating Committee report, 2010 election for the USENIX Board of Directors. ;login: the USENIX Association newsletter, 35(1): ??, February 2010. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/february-
2010-volume-35-number-1/nominating-
committee-report-2010-election-
usenix.

Escriva:2012:IHB

[EWS12] Robert Escriva, Bernard Wong, and Emin Gün Sirer. An introduction to HyperDex and the brave new world of high performance, scalable, consistent, fault-tolerant data stores. ;login: the USE-NIX Association newsletter, 37 (3):??, June 2012. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/june-
2012-volume-37-number-3/introduction-
hyperdex-and-brave-new-world-
high.

Farrow:2010:BMI

[Far10a]  Rik Farrow.  The Bar-

relfish multikernel: An interview with Timothy Roscoe. ;login: the USENIX Association newsletter, 35(2): ??, April 2010. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/april-
2010-volume-35-number-2/barrelfish-
multikernel-interview-timothy-
roscoe.

Farrow:2010:Ma

[Far10b] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 35(1):??, February 2010. CODEN LOGNEM. ISSN 1044-6397. URL https://www.

usenix.org/publications/login/

february-2010-volume-35-number-
1/musings.

Farrow:2010:Mb

[Far10c] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 35(2):??, April 2010. CODEN LOGNEM. ISSN 1044-6397. URL https://www.

usenix.org/publications/login/
april-2010-volume-35-number-
2/musings.

Farrow:2010:Mc

[Far10d] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 35(3):??, June 2010. CODEN LOGNEM. ISSN 1044-6397. URL https://www.

usenix.org/publications/login/
june-2010-volume-35-number-
3/musings.

Farrow:2010:Md

[Far10e] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 35(4):??, August 2010. CODEN LOGNEM. ISSN 1044-6397. URL https://www.

usenix.org/publications/login/
august-2010-volume-35-number-
4/musings.

Farrow:2010:Me

[Far10f] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 35(5):??, October 2010. CODEN LOGNEM. ISSN 1044-6397. URL https://www.

usenix.org/publications/login/
october-2010-volume-35-number-
5/musings.

Farrow:2010:Mf

[Far10g] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 35(6):??, December 2010. CODEN LOGNEM. ISSN 1044-6397. URL https://www.

usenix.org/publications/login/
december-2010-volume-35-number-
6/musings.

Farrow:2011:BSM

[Far11a] Rik Farrow. Backup strategies for molecular dynamics: An interview with Doug Hughes. ;login: the USENIX Association newsletter, 36(2):25– 28, April 2011. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/april-
2011-volume-36-number-2/backup-
strategies-molecular-dynamics-
interview-doug.

Farrow:2011:ISS

[Far11b] Rik Farrow. Interview with Stefan Savage: On the spam payment trail. ;login: the USE-NIX Association newsletter, 36 (4):??, August 2011. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/august-
2011-volume-36-number-4/interview-
stefan-savage-spam-payment-
trail.

Farrow:2011:Ma

[Far11c] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 36(1):??, February 2011. CODEN LOGNEM. ISSN 1044-6397. URL https://www.

usenix.org/publications/login/
february-2011-volume-36-number-
1/musings.

Farrow:2011:Mb

[Far11d] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 36(2):??, April 2011. CODEN LOGNEM. ISSN 1044-6397. URL https://www.

usenix.org/publications/login/
april-2011-volume-36-number-
2/musings.

Farrow:2011:Mc

[Far11e] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 36(3):??, June 2011. CODEN LOGNEM. ISSN 1044-6397. URL https://www.

usenix.org/publications/login/
june-2011-volume-36-number-
3/musings.

Farrow:2011:Md

[Far11f] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 36(4):??, August 2011. CODEN LOGNEM. ISSN 1044-6397. URL https://www.

usenix.org/publications/login/
august-2011-volume-36-number-
4/musings.

Farrow:2011:Me

[Far11g] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 36(5):??, October 2011. CODEN LOGNEM. ISSN 1044-6397. URL https://www.

usenix.org/publications/login/

october-2011-volume-36-number-
5/musings.

Farrow:2011:Mf

[Far11h] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 36(6):??, December 2011. CODEN LOGNEM. ISSN 1044-6397. URL https://www.

usenix.org/publications/login/
december-2011-volume-36-number-
6/musings.

Farrow:2012:JFM

[Far12a] Rik Farrow. Jacob Farmer on managing research data: An interview. ;login: the USE-NIX Association newsletter, 37 (5):??, October 2012. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/october-
2012-volume-37-number-5/jacob-
farmer-managing-research-
data-interview.

Farrow:2012:Ma

[Far12b] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 37(1):??, February 2012. CODEN LOGNEM. ISSN 1044-6397. URL https://www.

usenix.org/publications/login/
february-2012/musings.

Farrow:2012:Mb

[Far12c] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 37(2):??, April 2012. CODEN LOGNEM. ISSN 1044-6397. URL https://www.

usenix.org/publications/login/
april-2012/musings.

Farrow:2012:Mc

[Far12d] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 37(3):??, June 2012. CODEN LOGNEM. ISSN 1044-6397. URL https://www.

usenix.org/publications/login/
june-2012-volume-37-number-
3/musings.

Farrow:2012:Md

[Far12e] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 37(4):??, August 2012. CODEN LOGNEM. ISSN 1044-6397. URL https://www.

usenix.org/publications/login/
august-2012-volume-37-number-
4/musings.

Farrow:2012:Me

[Far12f] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 37(5):??, October 2012. CODEN LOGNEM. ISSN 1044-6397. URL https://www.

usenix.org/publications/login/
october-2012-volume-37-number-
5/musings.

Farrow:2012:Mf

[Far12g] Rik Farrow. Musings. ;login: the USENIX Association newsletter, 37(6):??, December 2012. CODEN LOGNEM. ISSN 1044-6397. URL https://www.

usenix.org/publications/login/
december-2012-volume-37-number-
6/musings.

Farrow:2012:NMC

[Far12h]  Rik Farrow.  Nathan Mil-

ford on Cassandra: An interview. ;login: the USE-NIX Association newsletter, 37 (3):??, June 2012. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/june-
2012-volume-37-number-3/nathan-
milford-cassandra-interview.

Farrow:2012:NHC

[Far12i]  Rik Farrow.  Netflix heads

into the clouds: Interview with Adrian Cockcroft. ;login: the USENIX Association newsletter, 37(1):??, February 2012. CODEN LOGNEM. ISSN 1044-6397. URL https://www.

usenix.org/publications/login/
february-2012/netflix-heads-
clouds-interview-adrian-cockcroft.

Ferrell:2010:DR

[Fer10a] Robert G. Ferrell. /dev/random. ;login: the USENIX Association newsletter, 35(6):??, December 2010. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/december-
2010-volume-35-number-6/devrandom.

Ferrell:2010:DRD

[Fer10b] Robert G. Ferrell. /dev/random: a debatably helpful guide to IT disaster planning. ;login: the USENIX Association newsletter, 35(4):??, August 2010. CODEN LOGNEM. ISSN 1044-6397. URL https://www.

usenix.org/publications/login/
august-2010-volume-35-number-
4/devrandom-debatably-helpful-
guide-it-disaster.

Ferrell:2010:DRA

[Fer10c] Robert G. Ferrell. /dev/random:

Airport security and other myths. ;login: the USENIX Association newsletter, 35(5): ??, October 2010. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/october-
2010-volume-35-number-5/devrandom-
airport-security-and-other-
myths.

Ferrell:2010:DRF

[Fer10d] Robert G. Ferrell. /dev/random:

Five common misconceptions about information security. ;login: the USENIX Association newsletter, 35(3): ??, June 2010. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/june-
2010-volume-35-number-3/devrandom-
five-common-misconceptions-
about.

Ferrell:2010:DRL

[Fer10e] Robert G. Ferrell. /dev/random:

Less successful network protocols. ;login: the USE-NIX Association newsletter, 35 (2):??, April 2010. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/april-
2010-volume-35-number-2/devrandom-
less-successful-network-protocols.

Ferrell:2010:DRP

[Fer10f] Robert G. Ferrell. /dev/random:

Project management — a primer. ;login: the USENIX Association newsletter, 35(1): ??, February 2010. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/february-
                                        2010-volume-35-number-1/devrandom-
                                        project-management%E2%80%
                                        94-primer.

Ferrell:2011:DRa

[Fer11a] Robert G. Ferrell. /dev/random. ;login: the USENIX Association newsletter, 36(1):??, February 2011. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/february-
                                        2011-volume-36-number-1/devrandom.

Ferrell:2011:DRb

[Fer11b] Robert G. Ferrell. /dev/random. ;login: the USENIX Association newsletter, 36(3): ??, June 2011. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/june-
                                        2011-volume-36-number-3/devrandom.

Ferrell:2011:DRc

[Fer11c] Robert G. Ferrell. /dev/random. ;login: the USENIX Association newsletter, 36(4): ??, August 2011. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/august-
                                        2011-volume-36-number-4/devrandom.

Ferrell:2011:DRd

[Fer11d] Robert G. Ferrell. /dev/random. ;login: the USENIX Association newsletter, 36(6):??, December 2011. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/december-
2011-volume-36-number-6/devrandom.

Ferrell:2011:DRP

[Fer11e] Robert G. Ferrell. /dev/random (parenthetically speaking). ;login: the USENIX Association newsletter, 36(2):??, April 2011. CODEN LOGNEM. ISSN 1044-6397. URL https://www.

usenix.org/publications/login/
april-2011-volume-36-number-
2/devrandom-parenthetically-
speaking.

Ferrell:2011:DRV

[Fer11f] Robert G. Ferrell. /dev/random:

Virtualization: a dark proverb. ;login: the USENIX Association newsletter, 36(5):??, October 2011. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/october-
2011-volume-36-number-5/devrandom-
virtualization-dark-proverb.

Ferrell:2012:DRa

[Fer12a] Robert G. Ferrell. /dev/random.

;login: the USENIX Association newsletter, 37(1):??, February 2012. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/february-
2012/devrandom.

Ferrell:2012:DRb

[Fer12b] Robert G. Ferrell. /dev/random.

;login: the USENIX Association newsletter, 37(3):

??, June 2012. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/june-
                                        2012-volume-37-number-3/devrandom.

Ferrell:2012:DRe

[Fer12c] Robert G. Ferrell. /dev/random. ;login: the USENIX Association newsletter, 37(5):??, October 2012. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/october-
                                        2012-volume-37-number-5/devrandom.

Ferrell:2012:DRf

[Fer12d] Robert G. Ferrell. /dev/random. ;login: the USENIX Association newsletter, 37(6):??, December 2012. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/december-
                                        2012-volume-37-number-6/devrandom.

Ferrell:2012:DRAa

[Fer12e] Robert G. Ferrell. /dev/random: Apps for saps. ;login: the USE-NIX Association newsletter, 37 (4):??, August 2012. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/august-
                                        2012-volume-37-number-4/devrandom-
                                        apps-saps.

Ferrell:2012:DRD

[Fer12f] Robert G. Ferrell. /dev/random: Dark rhetoric. ;login: the USE-NIX Association newsletter, 37 (2):??, April 2012. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/april-
2012/devrandom-dark-rhetoric.

Fiske:2011:MSA

[Fis11] Joshua Fiske. Making system administration easier by letting the machines do the hard work, or, becoming an agile sysadmin. ;login: the USE-NIX Association newsletter, 36 (1):??, February 2011. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/february-
2011-volume-36-number-1/making-
system-administration-easier-
letting.

Forno:2012:LWI

[For12] Richard F. Forno. Looking within to improve American cybersecurity. ;login: the USE-NIX Association newsletter, 37 (4):??, August 2012. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/august-
2012-volume-37-number-4/looking-
within-improve-american-cybersecurity.

Galvin:2010:PATf

[Gal10a] Peter Baer Galvin. Pete’s all things Sun: Comparing Solaris to RedHat Enterprise and AIX. ;login: the USENIX Association newsletter, 35(6): ??, December 2010. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/december-
2010-volume-35-number-6/petes-

                                        all-things-sun-comparing-
                                        solaris-redhat.

Galvin:2010:PATa

[Gal10b]  Peter Baer Galvin.  Pete’s

all things Sun: Crossbow. ;login: the USENIX Association newsletter, 35(1):??, February 2010. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/february-
                                        2010-volume-35-number-1/petes-
                                        all-things-sun-crossbow.

Galvin:2010:PATc

[Gal10c] Peter Baer Galvin. Pete’s all things Sun: Exadata V2 architecture and why it matters. ;login: the USENIX Association newsletter, 35(3): ??, June 2010. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/june-
                                        2010-volume-35-number-3/petes-
                                        all-things-sun-exadata-v2-
                                        architecture-and-why.

Galvin:2010:PATb

[Gal10d] Peter Baer Galvin. Pete’s all things Sun: Open source and free deduplication. ;login: the USENIX Association newsletter, 35(2):??, April 2010. CO-DEN LOGNEM. ISSN 1044-

                                        6397.  URL https://www.
                                        usenix.org/publications/login/
                                        april-2010-volume-35-number-
                                        2/petes-all-things-sun-open-
                                        source-and-free.

Galvin:2010:PATe

[Gal10e] Peter Baer Galvin. Pete’s all things Sun: The ‘Problem’ with NAS. ;login: the USENIX Association newsletter, 35(5): ??, October 2010. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/october-
2010-volume-35-number-5/petes-
all-things-sun-problem-nas.

Galvin:2010:PATd

[Gal10f] Peter Baer Galvin. Pete’s all things Sun: The status of Sun products. ;login: the USE-NIX Association newsletter, 35 (4):??, August 2010. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/august-
2010-volume-35-number-4/petes-
all-things-sun-status-sun-
products.

Galvin:2011:GATa

[Gal11a] Peter Baer Galvin. Galvin’s all things enterprise: The state of the Cloud. ;login: the USE-NIX Association newsletter, 36 (4):??, August 2011. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/august-
2011-volume-36-number-4/galvins-
all-things-enterprise-state-
cloud.

Galvin:2011:GATb

[Gal11b] Peter Baer Galvin. Galvin’s all things enterprise: The state of the Cloud, part

2. ;login: the USENIX Association newsletter, 36(5): ??, October 2011. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/october-
2011-volume-36-number-5/galvins-
all-things-enterprise-state-
cloud-part-2.

Galvin:2011:PATa

[Gal11c]  Peter Baer Galvin.  Pete’s

all things Sun: Comparing Solaris to RedHat Enterprise and AIX — virtualization features. ;login: the USENIX Association newsletter, 36(1): ??, February 2011. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/february-
2011-volume-36-number-1/petes-
all-things-sun-comparing-
solaris-redhat.

Galvin:2011:PATb

[Gal11d]  Peter Baer Galvin.  Pete’s

all things Sun: Solaris 11 Express. ;login: the USE-NIX Association newsletter, 36 (2):??, April 2011. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/april-
2011-volume-36-number-2/petes-
all-things-sun-solaris-11-
express.

Galvin:2012:SYC

[Gal12] Peter Baer Galvin. Should you care about Solaris 11? ;login: the USENIX Association newsletter, 37(2): ??, April 2012. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/april-
2012/should-you-care-about-
solaris-11.

Garfinkel:2012:PU

[Gar12] Simson L. Garfinkel. Programming Unicode. ;login: the USE-NIX Association newsletter, 37 (2):25–37, April 2012. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/april-
2012/programming-unicode.

Geer:2010:VC

[Gee10] Dan Geer, Jr. Vulnerable compliance. ;login: the USENIX Association newsletter, 35(6): ??, December 2010. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/december-
2010-volume-35-number-6/vulnerable-
compliance.

Geer:2012:FSC

[Gee12a]  Daniel E. Geer, Jr.  The

future of security: Criticality, rejectionists, risk tolerance. ;login: the USENIX Association newsletter, 37(4): ??, August 2012. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/august-
2012-volume-37-number-4/future-
security-criticality-rejectionists-
risk.

Geer:2012:FSN

[Gee12b] Daniel E. Geer, Jr. The future of security nominal delivery draft, Rocky Mountain Information Security Conference, 18 May 2012. ;login: the USE-NIX Association newsletter, 37 (4):??, August 2012. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/august-
2012-volume-37-number-4/future-
security-nominal-delivery-
draft-rocky.

Gutmann:2011:DUV

[Gut11] Peter Gutmann. Do users verify SSH keys? ;login: the USE-NIX Association newsletter, 36 (4):??, August 2011. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/august-
2011-volume-36-number-4/do-
users-verify-ssh-keys.

Hardavellas:2012:RFD

[Har12]  Nikos Hardavellas.  The

rise and fall of dark silicon. ;login: the USENIX Association newsletter, 37(2): ??, April 2012. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/april-
2012/rise-and-fall-dark-silicon.

Henderson:2012:N

[HD12] Casey Henderson and Anne Dickison. Notes. ;login: the USENIX Association newsletter, 37(4):??, August 2012. CO-DEN LOGNEM. ISSN 1044-

6397.  URL https://www.
usenix.org/publications/login/

august-2012-volume-37-number-
4/notes.

Hertel:2012:SMB

[Her12] Christopher R. Hertel. Server Message Block in the age of Microsoft glasnost. ;login: the USENIX Association newsletter, 37(1):??, February 2012. CODEN LOGNEM. ISSN 1044-6397. URL https://www.

usenix.org/publications/login/
february-2012/server-message-
block-age-microsoft-glasnost.

Hetzler:2011:SIS

[Het11]  Steven R. Hetzler.  Sys-

tem impacts of storage trends: Hard errors and testability. ;login: the USENIX Association newsletter, 36(3): ??, June 2011. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/june-
2011-volume-36-number-3/system-
impacts-storage-trends-hard-
errors-and.

Hicks:2010:OUC

[HFK^{+}10] Matthew Hicks, Murph Finnicum, Samuel T. King, Milo M. K. Martin, and Jonathan M. Smith. Overcoming an untrusted computing base: Detecting and removing malicious hardware automatically. ;login: the USENIX Association newsletter, 35(6): 31–41, December 2010. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/december-
2010-volume-35-number-6/overcoming-

                                        untrusted-computing-base-
                                        detecting.

Hindman:2011:MFR

[HKZ^{+}11] Benjamin Hindman, Andy Konwinski, Matei Zaharia, Ali Ghodsi, Anthony D. Joseph, Randy H. Katz, Scott Shenker, and Ion Stoica. Mesos: Flexible resource sharing for the Cloud. ;login: the USENIX Association newsletter, 36(4): ??, August 2011. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/august-
                                        2011-volume-36-number-4/mesos-
                                        flexible-resource-sharing-
                                        cloud.

Hughes:2012:WDCa

[HO12] Doug Hughes and Nathan Olla. When disasters collide: a manyfanged tale of woe and coincidence and patience and stress. ;login: the USE-NIX Association newsletter, 37 (3):??, June 2012. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/june-
                                        2012-volume-37-number-3/when-
                                        disasters-collide-many-fanged-
                                        tale-woe.

Hughes:2012:DIF

[Hug12a]  Doug Hughes.  Data in-

tegrity: Finding truth in a world of guesses and lies. ;login: the USENIX Association newsletter, 37(2): ??, April 2012. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/april-
2012/data-integrity-finding-
truth-world-guesses-and-lies.

Hughes:2012:WDCb

[Hug12b] Doug Hughes. When disasters collide: a many-fanged tale of woe, coincidence, patience, and stress, continued. ;login: the USENIX Association newsletter, 37(4): ??, August 2012. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/august-
2012-volume-37-number-4/when-
disasters-collide-many-fanged-
tale-woe.

Hawkey:2012:RAS

[HVB12] Kirstie Hawkey, Nicole Forsgren Velasquez, and Tamara Babaian. Report from the 5th ACM Symposium on Computer Human Interaction for Management of IT. ;login: the USE-NIX Association newsletter, 37 (2):??, April 2012. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/april-
2012/report-5th-acm-symposium-
computer-human-interaction-
management-it.

Josephson:2010:DFS

[JBFL10] William K. Josephson, Lars A.

Bongo, David Flynn, and Kai Li. DFS: a file system for virtualized flash storage. ;login: the USENIX Association newsletter, 35(2):

??, April 2010. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/april-
                                        2010-volume-35-number-2/dfs-
                                        file-system-virtualized-flash-
                                        storage.

Josephsen:2010:IQS

[Jos10a] Dave Josephsen. iVoyeur: a question of scale. ;login: the USENIX Association newsletter, 35(2):??, April 2010. CO-DEN LOGNEM. ISSN 1044-

                                        6397.  URL https://www.
                                        usenix.org/publications/login/
                                        april-2010-volume-35-number-
                                        2/ivoyeur-question-scale.

Josephsen:2010:IBH

[Jos10b]  Dave Josephsen.  iVoyeur:

Breaking up is hard to do. ;login: the USENIX Association newsletter, 35(1):??, February 2010. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/february-
                                        2010-volume-35-number-1/ivoyeur-
                                        breaking-hard-do.

Josephsen:2010:IGB

[Jos10c] Dave Josephsen. iVoyeur: Ganglia on the brain. ;login: the USENIX Association newsletter, 35(6):??, December 2010. CODEN LOGNEM. ISSN 1044-6397. URL https://www.

                                        usenix.org/publications/login/
                                        december-2010-volume-35-number-
                                        6/ivoyeur-ganglia-brain.

Josephsen:2010:IPPa

[Jos10d]  Dave Josephsen.  iVoyeur:

Pockets-o-packets, part 1. ;login: the USENIX Association newsletter, 35(3):??, June 2010. CODEN LOGNEM. ISSN 1044-6397. URL https://www.

usenix.org/publications/login/
june-2010-volume-35-number-
3/ivoyeur-pockets-o-packets-
part-1.

Josephsen:2010:IPPb

[Jos10e]  Dave Josephsen.  iVoyeur:

Pockets-o-packets, part 2. ;login: the USENIX Association newsletter, 35(4):??, August 2010. CODEN LOGNEM. ISSN 1044-6397. URL https://www.

usenix.org/publications/login/
august-2010-volume-35-number-
4/ivoyeur-pockets-o-packets-
part-2.

Josephsen:2010:IPPc

[Jos10f]  Dave Josephsen.  iVoyeur:

Pockets-o-packets, part 3. ;login: the USENIX Association newsletter, 35(5):??, October 2010. CODEN LOGNEM. ISSN 1044-6397. URL https://www.

usenix.org/publications/login/
october-2010-volume-35-number-
5/ivoyeur-pockets-o-packets-
part-3.

Josephsen:2011:IHG

[Jos11a] Dave Josephsen. iVoyeur: a handful of gems. ;login: the USENIX Association newsletter, 36(3):??, June 2011. CO-DEN LOGNEM. ISSN 1044-

                                        6397.  URL https://www.
                                        usenix.org/publications/login/
                                        june-2011-volume-36-number-
                                        3/ivoyeur-handful-gems.

Josephsen:2011:IAT

[Jos11b] Dave Josephsen. iVoyeur: All together now. ;login: the USE-NIX Association newsletter, 36 (2):??, April 2011. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/april-
                                        2011-volume-36-number-2/ivoyeur-
                                        all-together-now.

Josephsen:2011:ICGb

[Jos11c]  Dave Josephsen.  iVoyeur:

Changing the game. ;login: the USENIX Association newsletter, 36(6):??, December 2011. CODEN LOGNEM. ISSN 1044-6397. URL https://www.

                                        usenix.org/publications/login/
                                        december-2011-volume-36-number-
                                        6/ivoyeur-changing-game.

Josephsen:2011:ICGa

[Jos11d] Dave Josephsen. iVoyeur: Cloud gazing. ;login: the USENIX Association newsletter, 36(5): ??, October 2011. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/october-
                                        2011-volume-36-number-5/ivoyeur-
                                        cloud-gazing.

Josephsen:2011:IC

[Jos11e]  Dave Josephsen.  iVoyeur:

Crom. ;login: the USENIX Association newsletter, 36(4): ??, August 2011. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/august-
2011-volume-36-number-4/ivoyeur-
crom.

Josephsen:2011:IMG

[Jos11f]  Dave Josephsen.  iVoyeur:

More ganglia on the brain. ;login: the USENIX Association newsletter, 36(1):??, February 2011. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/february-
2011-volume-36-number-1/ivoyeur-
more-ganglia-brain.

Josephsen:2012:ICGa

[Jos12a]  Dave Josephsen.  iVoyeur:

Changing the game, part 2. ;login: the USENIX Association newsletter, 37(1):??, February 2012. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/february-
2012/ivoyeur-changing-game-
part-2.

Josephsen:2012:ICGb

[Jos12b]  Dave Josephsen.  iVoyeur:

Changing the game, part

3. ;login: the USENIX Association newsletter, 37(2): ??, April 2012. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/april-
2012/ivoyeur-changing-game-
part-3.

Josephsen:2012:ICGc

[Jos12c]  Dave Josephsen.  iVoyeur:

Changing the game, part

4. ;login: the USENIX Association newsletter, 37(3): ??, June 2012. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/june-
                                        2012-volume-37-number-3/ivoyeur-
                                        changing-game-part-4.

Josephsen:2012:IC

[Jos12d]  Dave Josephsen.  iVoyeur:

Crosstalk. ;login: the USE-NIX Association newsletter, 37 (5):??, October 2012. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/october-
                                        2012-volume-37-number-5/ivoyeur-
                                        crosstalk.

Josephsen:2012:INX

[Jos12e] Dave Josephsen. iVoyeur: Nagios XI. ;login: the USENIX Association newsletter, 37(6): ??, December 2012. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/december-
                                        2012-volume-37-number-6/ivoyeur-
                                        nagios-xi.

Josephsen:2012:IGF

[Jos12f] Dave Josephsen. iVoyeur: The gift of fire. ;login: the USE-NIX Association newsletter, 37 (4):??, August 2012. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/august-

2012-volume-37-number-4/ivoyeur-
gift-fire.

Jang:2012:RFU

[JWB12] Jiyong Jang, Maverick Woo, and David Brumley. ReDe-Bug: Finding unpatched code clones in entire OS distributions. ;login: the USENIX Association newsletter, 37(6): ??, December 2012. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/december-
2012-volume-37-number-6/redebug-
finding-unpatched-code-clones-
entire-os.

Kamath:2010:WL

[Kam10] Sean Kamath. Whither LISA?

;login: the USENIX Association newsletter, 35(1):??, February 2010. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/february-
2010-volume-35-number-1/whither-
lisa.

Knauerhase:2012:EPY

[KCT12] Rob Knauerhase, Romain Cledat, and Justin Teller. For extreme parallelism, your OS is sooooo last-millennium. ;login: the USENIX Association newsletter, 37(5):??, October 2012. CODEN LOGNEM. ISSN 1044-6397. URL https://www.

usenix.org/publications/login/
october-2012-volume-37-number-
5/extreme-parallelism-your-
os-sooooo-last.

Kendrick:2011:TTS

[Ken11] Stuart Kendrick. Testing the transport side of highly available hosts. ;login: the USE-NIX Association newsletter, 36 (3):??, June 2011. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/june-
                                        2011-volume-36-number-3/testing-
                                        transport-side-highly-available-
                                        hosts.

Kendrick:2012:TCN

[Ken12a] Stuart Kendrick. Tasting client/ network/server pie. ;login: the USENIX Association newsletter, 37(1):??, February 2012. CODEN LOGNEM. ISSN 1044-6397. URL https://www.

                                        usenix.org/publications/login/
                                        february-2012/tasting-clientnetworkserver-
                                        pie.

Kendrick:2012:WTU

[Ken12b] Stuart Kendrick. What takes us down? ;login: the USE-NIX Association newsletter, 37 (5):??, October 2012. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/october-
                                        2012-volume-37-number-5/what-
                                        takes-us-down.

Keromytis:2010:LVV

[Ker10]  Angelos D. Keromytis.  A

look at VoIP vulnerabilities. ;login: the USENIX Association newsletter, 35(1):??, February 2010. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/february-
2010-volume-35-number-1/look-
voip-vulnerabilities.

Khurana:2012:IHS

[Khu12] Amandeep Khurana. Introduction to HBase schema design. ;login: the USENIX Association newsletter, 37(5): ??, October 2012. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/october-
2012-volume-37-number-5/introduction-
hbase-schema-design.

Kirouac:2010:SEM

[Kir10] Brian Kirouac. Secure email for mobile devices. ;login: the USE-NIX Association newsletter, 35 (5):??, October 2010. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/october-
2010-volume-35-number-5/secure-
email-mobile-devices.

Kolstad:2010:UWW

[Kol10]  Rob Kolstad.  USA wins

World High School Programming Championships — snaps extended China winning streak. ;login: the USENIX Association newsletter, 35(5):??, October 2010. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/october-
2010-volume-35-number-5/usa-
wins-world-high-school-programming.

Kemerlis:2012:KLK

[KPAK12] Vasileios Kemerlis, Georgios Portokalidis, Elias Athanasopoulos, and Angelos D. Keromytis. kGuard: Lightweight kernel protection. ;login: the USENIX Association newsletter, 37(6):??, December 2012. CODEN LOGNEM. ISSN 1044-6397. URL https://www.

                                        usenix.org/publications/login/
                                        december-2012-volume-37-number-
                                        6/kguard-lightweight-kernel-
                                        protection.

Katz:2010:MLL

[KS10] Lou Katz and Peter H. Salus. In memoriam: Lewis A. Law, 1932–2010. ;login: the USE-NIX Association newsletter, 35 (2):??, April 2010. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/april-
                                        2010-volume-35-number-2/memoriam-
                                        lewis-law-1932%E2%80%932010.

Langley:2011:BBH

[Lan11]  Adam Langley.  Beyond

the basics of HTTPS serving. ;login: the USENIX Association newsletter, 36(6): ??, December 2011. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/december-
                                        2011-volume-36-number-6/beyond-
                                        basics-https-serving.

Leidner:2010:BIH

[LB10] Jochen L. Leidner and Gary Berosik. Building and installing a Hadoop/MapReduce cluster from commodity components: a case study. ;login: the USE-NIX Association newsletter, 35 (1):??, February 2010. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/february-
2010-volume-35-number-1/building-
and-installing-hadoopmapreduce-
cluster.

Lamourine:2012:BR

[LCS^{+}12] Mark Lamourine, Brandon Ching, Peter H. Salus, Jeff Berg, Evan Teran, and Rik Farrow. Book reviews. ;login: the USE-NIX Association newsletter, 37 (3):??, June 2012. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/june-
2012-volume-37-number-3/book-
reviews.

Limoncelli:2011:LTD

[LH11] Thomas A. Limoncelli and Doug Hughes. LISA ’11: Theme — ‘DevOps: New challenges and proven values’. ;login: the USE-NIX Association newsletter, 36 (4):??, August 2011. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/august-
2011-volume-36-number-4/lisa-
11-theme%E2%80%94devops-new-
challenges-proven-values.

Limoncelli:2010:SAP

[Lim10]  Thomas A. Limoncelli.  A

system administration parable: The waitress and the water glass. ;login: the USENIX Association newsletter, 35(5):

??, October 2010. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/october-
                                        2010-volume-35-number-5/system-
                                        administration-parable-waitress-
                                        and-water.

Li:2010:SPS

[LRA10] Bo Li, Elena Reshetova, and Tuomas Aura. Symbian OS platform security model. ;login: the USENIX Association newsletter, 35(4): ??, August 2010. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/august-
                                        2010-volume-35-number-4/symbian-
                                        os-platform-security-model.

Lueninghoener:2011:GSC

[Lue11] Cory Lueninghoener. Getting started with configuration management. ;login: the USE-NIX Association newsletter, 36 (2):??, April 2011. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/april-
                                        2011-volume-36-number-2/getting-
                                        started-configuration-management.

McDonald:2012:N

[McD12]  Alex McDonald.  NFSv4.

;login: the USENIX Association newsletter, 37(1):??, February 2012. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/february-
                                        2012/nfsv4.

McKee:2010:MHE

[McK10] Troy McKee. Migrating from hosted Exchange. ;login: the USENIX Association newsletter, 35(5):??, October 2010. CO-DEN LOGNEM. ISSN 1044-

6397.  URL https://www.
usenix.org/publications/login/
october-2010-volume-35-number-
5/migrating-hosted-exchange.

Murray:2011:SCP

[MH11] Derek G. Murray and Steven Hand. Skywriting on CIEL: Programming the data center. ;login: the USENIX Association newsletter, 36(2): ??, April 2011. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/april-
2011-volume-36-number-2/skywriting-
ciel-programming-data-center.

Maltzahn:2010:CSA

[MMEK^{+}10] Carlos Maltzahn, Esteban Molina-Estolano, Amandeep Khurana, Alex J. Nelson, Scott A. Brandt, and Sage Weil. Ceph as a scalable alternative to the Hadoop distributed file system. ;login: the USENIX Association newsletter, 35(4): ??, August 2010. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/august-
2010-volume-35-number-4/ceph-
scalable-alternative-hadoop-
distributed-file.

Madhyastha:2012:SIP

[MMP^{+}12] Harsha V. Madhyastha, John C. Mccullough, George Porter, Rishi Kapoor, Stefan Savage, Alex C. Snoeren, and Amin Vahdat. scc: Informed provisioning of storage for cluster applications. ;login: the USE-NIX Association newsletter, 37 (3):??, June 2012. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/june-
2012-volume-37-number-3/scc-
informed-provisioning-storage-
cluster.

Moran:2011:UAP

[Mor11] Ned Moran. Understanding advanced persistent threats: a case study. ;login: the USE-NIX Association newsletter, 36 (4):??, August 2011. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/august-
2011-volume-36-number-4/understanding-
advanced-persistent-threats-
case.

Mudge:2011:NAC

[Mud11] Raphael Mudge. Network attack collaboration: Sharing the Shell. ;login: the USENIX Association newsletter, 36(4): ??, August 2011. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/august-
2011-volume-36-number-4/network-
attack-collaboration-sharing-
shell.

Mannan:2012:PBMb

[MV12a] Mohammad Mannan and P. C.

Van Oorschot. Passwords for both mobile and desktop computers appendix. ;login: the USENIX Association newsletter, 37(4):??, August 2012. CO-DEN LOGNEM. ISSN 1044-

6397.  URL https://www.
usenix.org/publications/login/
august-2012-volume-37-number-
4/passwords-both-mobile-and-
desktop-computers.

Mannan:2012:PBMa

[MV12b] Mohammad Mannan and P. C.

Van Oorschot. Passwords for both mobile and desktop computers: ObPwd for Firefox and Android. ;login: the USE-NIX Association newsletter, 37 (4):??, August 2012. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/august-
2012-volume-37-number-4/passwords-
both-mobile-and-desktop-computers-
obpwd.

Meyer:2011:NMV

[MWHW11] Dutch T. Meyer, Jake Wires, Norman C. Hutchinson, and Andrew Warfield. Namespace management in virtual desktops. ;login: the USENIX Association newsletter, 36(1): ??, February 2011. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/february-
2011-volume-36-number-1/namespace-
management-virtual-desktops.

Newton:2011:RWS

[NPFS11] Andrew Newton, Dave Piscitello, Benedetto Fiorelli, and Steve Sheng. A RESTful Web service for Internet name and address directory services. ;login: the USENIX Association newsletter, 36(5): ??, October 2011. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/october-
                                        2011-volume-36-number-5/restful-
                                        web-service-internet-name-
                                        and-address.

Oetiker:2012:HPA

[Oet12]  Tobi Oetiker.  How Perl

added Unicode support 10 years ago without you noticing it. ;login: the USE-NIX Association newsletter, 37 (2):??, April 2012. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/april-
                                        2012/how-perl-added-unicode-
                                        support-10-years-ago-without-
                                        you-noticing-it.

Piscitello:2010:SSI

[Pis10] David Piscitello. SSR: Securing the Internet’s identifier systems. ;login: the USENIX Association newsletter, 35(1): ??, February 2010. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/february-
                                        2010-volume-35-number-1/ssr-
                                        securing-internets-identifier-
                                        systems.

Piatek:2010:IPR

[PK10] Michael Piatek and Arvind Krishnamurthy. Improving the performance and robustness of P2P live streaming with contracts. ;login: the USENIX Association newsletter, 35(4): ??, August 2010. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/august-
2010-volume-35-number-4/improving-
performance-and-robustness-
p2p-live.

Polisher:2012:SPP

[Pol12] Charles Polisher. The summer of popping Power supplies. ;login: the USENIX Association newsletter, 37(5): ??, October 2012. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/october-
2012-volume-37-number-5/summer-
popping-power-supplies.

Rajab:2010:NEW

[RBM^{+}10] Moheeb Abu Rajab, Lucas Ballard, Panayiotis Mavrommatis, Niels Provos, and Xin Zhao. The Nocebo Effect on the Web: An analysis of fake anti-virus distribution. ;login: the USENIX Association newsletter, 35(6): ??, December 2010. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/december-
2010-volume-35-number-6/nocebo-
effect-web-analysis-fake-
anti-virus.

Revelle:2011:HVM

[Rev11] Don Revelle. Hypervisors and virtual machines: Implementation insights on the x86 architecture. ;login: the USENIX Association newsletter, 36(5): 17–22, October 2011. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/october-
                                        2011-volume-36-number-5/hypervisors-
                                        and-virtual-machines-implementation.

Roesner:2012:SBP

[RRKW12] Franziska Roesner, Christopher Rovillos, Tadayoshi Kohno, and David Wetherall. ShareMeNot: Balancing privacy and functionality of third-party social widgets. ;login: the USENIX Association newsletter, 37(4): ??, August 2012. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/august-
                                        2012-volume-37-number-4/sharemenot-
                                        balancing-privacy-and-functionality.

Ryanczak:2010:ITW

[Rya10] Matt Ryanczak. IPv6 transit: What you need to know. ;login: the USENIX Association newsletter, 35(5): ??, October 2010. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/october-
                                        2010-volume-35-number-5/ipv6-
                                        transit-what-you-need-know.

Sangwine:2010:WC

[San10] Jim Sangwine. 3D Web contenders. ;login: the USE-NIX Association newsletter, 35 (3):??, June 2010. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/june-
2010-volume-35-number-3/3d-
web-contenders.

Schaumann:2010:TSA

[Sch10]  Jan Schaumann.  Teaching

system administration in the Cloud. ;login: the USENIX Association newsletter, 35(5): ??, October 2010. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/october-
2010-volume-35-number-5/teaching-
system-administration-cloud.

Schaumann:2011:HUA

[Sch11a] Jan Schaumann. Of headless user accounts and restricted shells. ;login: the USENIX Association newsletter, 36(6):38– 44, December 2011. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/december-
2011-volume-36-number-6/headless-
user-accounts-and-restricted-
shells.

Schwartz:2011:DUC

[Sch11b]  Edward J. Schwartz.  The

danger of unrandomized code. ;login: the USENIX Association newsletter, 36(6):??, December 2011. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/december-
2011-volume-36-number-6/danger-
unrandomized-code.

Shekhar:2012:ASS

[SDW12] Shashi Shekhar, Michael Dietz, and Dan S. Wallach. AdSplit: Separating smartphone advertising from applications. ;login: the USENIX Association newsletter, 37(6): ??, December 2012. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/december-
                                        2012-volume-37-number-6/adsplit-
                                        separating-smartphone-advertising.

Seely:2010:BVD

[See10] Andrew Seely. Building a virtual DNS appliance using Solaris 10, BIND, and VMware. ;login: the USENIX Association newsletter, 35(3):27– 34, June 2010. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/june-
                                        2010-volume-35-number-3/building-
                                        virtual-dns-appliance-using-
                                        solaris-10-bind.

Sherwood:2011:SUY

[She11] Rob Sherwood. Safely using your production network as a testbed. ;login: the USENIX Association newsletter, 36(1): ??, February 2011. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/february-
                                        2011-volume-36-number-1/safely-
                                        using-your-production-network-
                                        testbed.

Shvachko:2010:HSL

[Shv10] Konstantin V. Shvachko. HDFS scalability: The limits to growth. ;login: the USE-NIX Association newsletter, 35 (2):??, April 2010. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/april-
2010-volume-35-number-2/hdfs-
scalability-limits-growth.

Shvachko:2011:AHS

[Shv11] Konstantin V. Shvachko. Apache Hadoop: The scalability update. ;login: the USENIX Association newsletter, 36(3): ??, June 2011. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/june-
2011-volume-36-number-3/apache-
hadoop-scalability-update.

Simmons:2010:LBJ

[Sim10] Matt Simmons. LISA: Beginning the journey. ;login: the USENIX Association newsletter, 35(3):??, June 2010. CO-DEN LOGNEM. ISSN 1044-

6397.  URL https://www.
usenix.org/publications/login/
june-2010-volume-35-number-
3/lisa-beginning-journey.

Slik:2012:IC

[Sli12] David Slik. An introduction to CDMI. ;login: the USE-NIX Association newsletter, 37 (3):??, June 2012. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/june-
2012-volume-37-number-3/introduction-
cdmi.

Sassaman:2011:HPN

[SPBS11] Len Sassaman, Meredith L. Patterson, Sergey Bratus, and Anna Shubina. The halting Problems of network stack insecurity. ;login: the USENIX Association newsletter, 36(6): ??, December 2011. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/december-
                                        2011-volume-36-number-6/halting-
                                        problems-network-stack-insecurity.

Tanenbaum:2010:MSR

[TAB^{+}10] Andrew Tanenbaum, Raja Appuswamy, Herbert Bos, Lorenzo Cavallaro, Cristiano Giuffrida, Tomás Hrubý, Jorrit Herder, Erik van der Kouwe, and David van Moolenbroek. MINIX 3: Status report and current research. ;login: the USENIX Association newsletter, 35(3): 7–13, June 2010. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/june-
                                        2010-volume-35-number-3/minix-
                                        3-status-report-and-current-
                                        research.

Tange:2011:GPC

[Tan11a]  Ole Tange.  GNU Paral-

lel: The command-line power tool. ;login: the USENIX Association newsletter, 36(1): ??, February 2011. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/february-
                                        2011-volume-36-number-1/gnu-

parallel-command-line-power-
tool.

Tange:2011:GSC

[Tan11b] Ole Tange. GNU SQL: a command line tool for accessing different databases using DBURLs. ;login: the USE-NIX Association newsletter, 36 (2):??, April 2011. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/april-
2011-volume-36-number-2/gnu-
sql-command-line-tool-accessing-
different.

Ur:2012:HUC

[UKK^{+}12] Blase Ur, Patrick Gage Kelley, Saranga Komanduri, Joel Lee, Michael Maass, Michelle L. Mazurek, Timothy Passaro, Richard Shay, Timothy Vidas, Lujo Bauer, Nicolas Christin, Lorrie Faith Cranor, Serge Egelman, and Julio López. Helping users create better passwords. ;login: the USENIX Association newsletter, 37(6): ??, December 2012. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/december-
2012-volume-37-number-6/helping-
users-create-better-passwords.

vanDrunen:2010:PRD

[vD10a] Rudi van Drunen. Peculiarities of radio devices. ;login: the USENIX Association newsletter, 35(1):64–72, February 2010. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/february-
                                        2010-volume-35-number-1/peculiarities-
                                        radio-devices.

vanDrunen:2010:SES

[vD10b] Rudi van Drunen. Small embedded systems. ;login: the USE-NIX Association newsletter, 35 (3):??, June 2010. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/june-
                                        2010-volume-35-number-3/small-
                                        embedded-systems.

Vixie:2011:OUS

[Vix11] Paul Vixie. Other uses for Secure DNS. ;login: the USE-NIX Association newsletter, 36 (5):??, October 2011. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/october-
                                        2011-volume-36-number-5/other-
                                        uses-secure-dns.

Watson:2010:ICP

[WALK10] Robert N. M. Watson, Jonathan Anderson, Ben Laurie, and Kris Kennaway. Introducing Capsicum: Practical capabilities for UNIX. ;login: the USENIX Association newsletter, 35(6): ??, December 2010. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/december-
                                        2010-volume-35-number-6/introducing-
                                        capsicum-practical-capabilities-
                                        unix.

Wanguhgu:2012:VEN

[Wan12] Kamau Wanguhgu. VXLAN:

Extending networking to fit the Cloud. ;login: the USENIX Association newsletter, 37(5): ??, October 2012. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/october-
2012-volume-37-number-5/vxlan-
extending-networking-fit-
cloud.

Winstein:2012:MSA

[WB12] Keith Winstein and Hari Balakrishnan. Mosh: a state-of-the-art good old-fashioned mobile shell. ;login: the USE-NIX Association newsletter, 37 (4):??, August 2012. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/august-
2012-volume-37-number-4/mosh-
state-art-good-old-fashioned-
mobile-shell.

Ching:2010:BRa

[wBC10] Elizabeth Zwicky with Brandon Ching. Book reviews. ;login: the USENIX Association newsletter, 35(2): ??, April 2010. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/april-
2010-volume-35-number-2/book-
reviews.

Winter:2012:GFC

[WC12] Philipp Winter and Jedidiah Crandall. The Great Firewall of China: How it blocks Tor and why it is hard to pinpoint. ;login: the USENIX Association newsletter, 37(6): ??, December 2012. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/december-
2012-volume-37-number-6/great-
firewall-china-how-it-blocks-
tor-and-why.

Stover:2012:BR

[wSS12] Elizabeth Zwicky. with Sam Stover. Book reviews. ;login: the USE-NIX Association newsletter, 37 (1):??, February 2012. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/february-
2012/book-reviews.

Wires:2010:BBF

[WW10] Jake Wires and Andrew Warfield. Beyond blocks and files. ;login: the USENIX Association newsletter, 35(2):??, April 2010. CODEN LOGNEM. ISSN 1044-6397. URL https://www.

usenix.org/publications/login/
april-2010-volume-35-number-
2/beyond-blocks-and-files.

Young:2010:TOVa

[You10a] Ellie Young. Thanks to our volunteers. ;login: the USE-NIX Association newsletter, 35 (1):??, February 2010. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/february-
2010-volume-35-number-1/thanks-
our-volunteers.

Young:2010:TOVb

[You10b] Ellie Young. Thanks to our volunteers. ;login: the USE-NIX Association newsletter, 35 (6):??, December 2010. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/december-
2010-volume-35-number-6/thanks-
our-volunteers.

etal:2010:BR

[Z^{+}10] Elizabeth Zwicky et al. Book reviews. ;login: the USENIX Association newsletter, 35(1): ??, February 2010. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/february-
2010-volume-35-number-1/book-
reviews.

Zaharia:2012:FIA

[ZCD^{+}12] Matei Zaharia, Mosharaf Chowdhury, Tathagata Das, Ankur Dave, Justin Ma, Murphy Mccauley, Michael J. Franklin, Scott Shenker, and Ion Stoica. Fast and interactive analytics over Hadoop data with Spark. ;login: the USENIX Association newsletter, 37(4): ??, August 2012. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/august-
2012-volume-37-number-4/fast-
and-interactive-analytics-
over-hadoop-data.

Zwicky:2010:BRa

[ZCS10a] Elizabeth Zwicky, Brandon Ching, and Sam Stover. Book reviews. ;login: the USENIX Association newsletter, 35(4): ??, August 2010. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/august-
2010-volume-35-number-4/book-
reviews.

Zwicky:2010:BRb

[ZCS10b] Elizabeth Zwicky, Brandon Ching, and Sam Stover. Book reviews. ;login: the USENIX Association newsletter, 35(5): ??, October 2010. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/october-
2010-volume-35-number-5/book-
reviews.

Zwicky:2011:BRe

[ZDSG11] Elizabeth Zwicky, Trey Darley, Sam Stover, and Peter Baer Galvin. Book reviews. ;login: the USENIX Association newsletter, 36(5): ??, October 2011. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/october-
2011-volume-36-number-5/book-
reviews.

Zwicky:2011:BRd

[ZF11] Elizabeth Zwicky and Rik Farrow. Book reviews. ;login: the USENIX Association newsletter, 36(4):??, August 2011. CODEN LOGNEM. ISSN 1044-6397. URL https://www.

usenix.org/publications/login/
august-2011-volume-36-number-
4/book-reviews.

Zwicky:2012:BRd

[ZF12] Elizabeth Zwicky and Rik Farrow. Book reviews. ;login: the USENIX Association newsletter, 37(6):??, December 2012. CODEN LOGNEM. ISSN 1044-6397. URL https://www.

usenix.org/publications/login/
december-2012-volume-37-number-
6/book-reviews.

Zwicky:2010:BRc

[ZFS10] Elizabeth Zwicky, Rik Farrow, and Sam Stover. Book reviews. ;login: the USENIX Association newsletter, 35(6): ??, December 2010. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/december-
2010-volume-35-number-6/book-
reviews.

Zwicky:2011:BRa

[ZFS11] Elizabeth Zwicky, Rik Farrow, and Sam Stover. Book reviews. ;login: the USENIX Association newsletter, 36(1): ??, February 2011. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/february-
2011-volume-36-number-1/book-
reviews.

Zwicky:2012:BRb

[ZGSL12] Elizabeth Zwicky, With Peter Gutmann, Sam Stover, and Mark Lamourine. Book reviews. ;login: the USENIX Association newsletter, 37(4): ??, August 2012. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/august-
                                        2012-volume-37-number-4/book-
                                        reviews.

Zwicky:2012:BRc

[ZL12] Elizabeth Zwicky and Mark Lamourine. Book reviews. ;login: the USENIX Association newsletter, 37(5):??, October 2012. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/october-
                                        2012-volume-37-number-5/book-
                                        reviews.

Zwicky:2012:BRa

[ZLDC12] Elizabeth Zwicky, Mark Lamourine, Trey Darley, and Brandon Ching. Book reviews. ;login: the USENIX Association newsletter, 37(2): ??, April 2012. CODEN LOGNEM. ISSN 1044-6397.

                                        URL https://www.usenix.
                                        org/publications/login/april-
                                        2012/book-reviews.

Zwicky:2011:BRf

[ZS11] Elizabeth Zwicky and Sam Stover. Book reviews. ;login: the USENIX Association newsletter, 36(6):??, December 2011. CODEN LOGNEM. ISSN 1044-6397. URL https://www.

                                        usenix.org/publications/login/
                                        december-2011-volume-36-number-
                                        6/book-reviews.

Zwicky:2011:BRb

[ZSTF11] Elizabeth Zwicky, Sam Stover, Evan Teran, and Rik Farrow.

Book reviews. ;login: the USE-NIX Association newsletter, 36 (2):??, April 2011. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/april-
2011-volume-36-number-2/book-
reviews.

Zwicky:2011:BRc

[ZTM11] Elizabeth Zwicky, Evan Teran, and Ryan MacArthur. Book reviews. ;login: the USE-NIX Association newsletter, 36 (3):??, June 2011. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/june-
2011-volume-36-number-3/book-
reviews.

Zadok:2010:CSF

[ZTS10] Erez Zadok, Vasily Tarasov, and Priya Sehgal. The case for specialized file systems, or, fighting file system obesity. ;login: the USENIX Association newsletter, 35(1): ??, February 2010. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/february-
2010-volume-35-number-1/case-
specialized-file-systems-
or-fighting-file.

Zwicky:2010:BFI

[Zwi10] Elizabeth Zwicky. Brute force and ignorance. ;login: the USE-NIX Association newsletter, 35 (2):??, April 2010. CODEN LOGNEM. ISSN 1044-6397.

URL https://www.usenix.
org/publications/login/april-

                                        2010-volume-35-number-2/brute-
                                        force-and-ignorance.